connect FINANCE BCOR 340: Spring 2019 MWF Section 2 omework 1 (Financial Statements Question 5 (of 12) 5. 1000 points value Problem 3-5 Calculating Leverage Ratios [LO 2 Allen, Inc., has a total debt ratio of .34 What is its debt-equity ratio? (Do not round intermediate calculations. Round your answer to 2 decimal places (o.g.. 32.16).) Debl-equity ratio Requirement 2: What is its equity multiplier? (Do not round intermediate calculations. Round your answer to 2 decimal places (e.g. 32.16)) Equity multiplier Hints References eBook & Resources Hint#1 esc Search or type URL
Answer #1

Total Debt/Total Assets=0.34

D/(D+E)=0.34
=>1/(1+E/D)=0.34
=>1/0.34-1=E/D
=>D/E=1/(1/0.34-1)
=>D/E=0.515151515
=>Debt/Equity=0.515151515

Equity multiplier=1+D/E=1+0.515151515=1.515151515